Fermi National Laboratory


Accelerator Update for May 13, 2005 - May 16, 2005

Friday May 13
The day shift began with Operations monitoring the Recycler stash, with Tevatron (TeV) experts conducting studies, with the Antiproton Source (Pbar) stacking, and with NuMI and the Meson MCenter experiments taking beam. MiniBooNE was off because of an LCW problem.

A MiniBooNE expert reported that the problem with the LCW for the horn was only an instrumentation problem and the horn could be turned back on. MiniBooNE resumed taking beam at 9:30 AM.

At 10:38 AM, a Booster kicker (MKS07) began causing trouble. Experts found a bad cable. They bypassed the station, repaired the cable, and had the kicker back on by 11:14 AM.

At 4:28 PM, two Booster vacuum valves closed, and a power supply tripped off (GMPS), a chopper (B:CHOP), and various other devices tripped off. Everything reset okay.

The TeV studies ended around 5 PM, but other TeV experts had to come in to work on problems with the abort gap.

At 5:18 PM, a Booster power supply (EAPS) tripped off for the first of three times in eleven minutes. It reset okay each time.

Operators made the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls ready for access at 6:41 PM.

Operations put the TeV into shot setup at 9:37 PM.

Operations established combination store 4144 at 11:49 PM with an initial luminosity of 123.09E30. Pbar resumed stacking.

Operation established beam to NuMI and MCenter by 11:56 PM.

Saturday May 14
The shift began with Operations monitoring store 4144 and a small stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:57 AM, Operations reported that the EAPS tripped off three times. On the last trip it wouldn't reset. Operators investigated. They were able to get the supply back on after they disconnected a Booster RF station (BRF4) from the system at 1:21 AM.

Operators managed to bring BRF4 back on line at 1:37 AM.

At 1:57 AM, the I- Source tripped off and wouldn't reset. Operators accessed the dome and then called in experts to help. The experts had the source back up by 3:14 AM.

At 4:34 AM, a Linac quadrupole power supply (Q107) began suffering voltage drops. Operators switched its control card at 4:50 AM.

At 5:31 AM, Pbar lost its beam permit, which caused NuMI to lose its permit also.

At 7:01 AM, BRF16 tripped off for the first of three times in eight minutes. It reset okay both times.

At 7:51 AM, BRF16 tripped off. The station resets, but won't stay on. Operators bypassed the station. RF technicians will repair the station.

Operations put the Recycler into shot setup to stash antiprotons at 8:54 AM. The transfers were completed by 9:39 AM. Pbar resumed stacking and NuMI and MCenter resumed taking beam by 10:07 AM.

At 10:27 AM, the Booster EAPS tripped off. An expert was here and investigated. He had the supply back on three minutes later.

The EAPS tripped off at 10:46 AM, and then three more times over the next fifteen minutes. It reset each time.

At 12:22 PM, BRF16 tripped off. It reset okay.

The Recycler began stashing antiprotons at 3:36 PM. The stash completed at 4:49 PM. Pbar resumed stacking and NuMI and MCenter resumed taking beam.

Sunday May 15
The midnight shift began with Operations monitoring store 4144 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 4:29 AM, the TeV suffered a sector A4 quench. Operators and cryo system technicians investigated and discovered that a wet engine had failed. Wet engine techs were called in.

Operators prepared CDF and D0 for collision hall access around 8 AM.

Cryo system techs reported at 8:28 AM that the A4 wet engine had a sheared bolt.

At 9:41 AM, Operations reported the I- Source started sparking. Operators investigated and found a failed power supply. They had beam back by 10:05 AM.

Cryo techs reported they had A4 wet engine repaired by 10:50 AM.

At 12:03 PM, the I- Source went down again. Operators reset a NIM crate.

Operations put the TeV into shot setup at 12:30 PM.

At 2:11 PM, the TeV suffered a sector E4 quench. The quench occurred during the second transfer from the Recycler. A timing event was left active that caused an injection kicker to fire.

NuMI resumed taking beam at 3:01 PM.

At 4:56 PM, Operators reported that there were no readings from the TeV dampers. Operators investigated and found a failed crate power supply at D2. They replaced the power supply at 5:25 PM and began a file restore.

Operations established combination store 4147 at 6:39 PM with an initial luminosity of 105.9E30. Pbar resumed stacking.

At 6:44 PM, the EAPS tripped off for the first of two times in ten minutes. It reset okay.

At 7:26 PM, the NuMI beam permit tripped off when a power supply (HV101) lost its ramp. Operators called in power supply techs.

At 8:40 PM, the EAPS tripped off for the first of two times in ten minutes. It reset okay.

The NuMI power supply work finished and NuMI beam resumed at 9:02 PM.

A Pbar expert began a two-hour Debuncher study around 9:06 PM

At 9:19 PM, BRF16 tripped off for the first of two times in eleven minutes. It reset okay.

Monday May 16
The midnight shift began with Operations monitoring store 4147 and the stash, with Pbar stacking, and with MiniBooNE , NuMI, and MCenter taking beam.

Recycler began stashing antiprotons at 3:56 AM.

At 5:28 AM, the Booster chopper tripped off. Operations called in experts. At 6:14 AM, they reported there was no power to the chopper. At 8 AM, they reported that they had replaced the power supply, but the chopped required sixty to ninety minutes to condition before they could find out if there was anything else wrong. This halts all beam.

The Plan for Monday
The plan for today is to repair the chopper and then resume stacking. Hopefully we will be able to have enough of a stack to establish store this evening.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/16/2005   email Fermilab